OpenStudy (anonymous):

Solve x2 - 12x + 5 = 0 by completing the square. Which equation is used in the process? (x - 6)2 = 31 , (x - 6)2 = 26 , (x - 6)2 = 19

OpenStudy (anonymous):

first one

OpenStudy (mertsj):

Subtract 5 from both sides. x^2-12x = -5 Take 1/2 of 12 which is 6 and square it and add that to both sides. x^2-12x+36=-5+36

OpenStudy (mertsj):

(x-6)^2=31 Take the square root of both sides. x-6=+-sqrt31 x=6+-sqrt31

OpenStudy (mertsj):

So first choice